# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v3

@@ 3 Learning with arbitrary effective costs
In this section, we consider the effective cost of agent $i$, i.e., $z_{i}(s_{i},s_{-i})$, to be an arbitrary function of $s=(s_{i},s_{-i})$. We know that an alternative interpretation of NE (Definition˜1) is a strategy profile $(s_{i}^{*},s_{-i}^{*})$ where every agent’s best response to the strategies of the other players is its own strategy in that profile (see  (Maschler et al., 2020, e.g.)), i.e., $s_{i}^{*}\in\operatorname*{\arg\!\max}_{s_{i}\in S_{i}}\ u_{i}(w,s_{i},s_{-i}^{*}),\ \forall i\in N$. Hence, an algorithm that simultaneously updates all agents’ strategies with the best responses to the current strategies of the other players is called a best response dynamics of a strategic form game (see  (Fudenberg, 1991, e.g.) for a detailed description). In our problem, this approach cannot be directly employed since the center also needs to learn and update the model parameters $w$ as the agents choose their data contributions. We, therefore, propose a mechanism for federated learning that simultaneously updates both the agents’ strategies and the center’s choice of $w$. Federated learning protocols iteratively use local gradients to update a global model, the mechanism we describe below captures strategic behaviour at every iteration by allowing agents to update their strategy (data set sizes used for gradient computation) at every realisation of the global model based on Equation˜1
@@
In this mechanism, each agent $i$ starts with some initial choices of $s_{i}^{0}$ and shares that with the center. [^1] Note that, only the number of data points, $s_{i}$ is shared with the center and not the data, which is consistent with the principle of federated learning. The center also starts with a $w^{0}$ and broadcasts that and the entire initial data-contribution vector $s^{0}$ to all the agents. In every subsequent iteration $t$, each agent $i$ locally computes two quantities: (i) its updated contribution $s_{i}^{t+1}$ by taking one gradient ascent step w.r.t. its own contribution $s_{i}$, and (ii) the local gradient of agent $i$’s valuation component of the social welfare w.r.t. $w$. Both are evaluated at the current values of $w^{t}$ and $s^{t}$ and sent back to the center. The center calculates an updated $w^{t+1}$ that averages (in spirit of the FedAvg algorithm (McMahan et al., 2017)) all the local gradients sent by the agents. The center then shares $w^{t+1}$ and $s^{t+1}$ with all the agents. Formally, the updates are given as follows.
$s^{t+1}$ $=s^{t}+\gamma g(w^{t},s^{t},\mu^{t}),$ (3)
$w^{t+1}$ $=w^{t}+\eta\tilde{g}(w^{t},s^{t}),$ (4)
where the function $\tilde{g}(w^{t},s^{t})=\frac{1}{n}\sum_{i\in N}\nabla_{w}v_{i}(w^{t},s^{t})$ and $g$ is defined as
$[g(w^{t},s^{t},\mu^{t})]_{i}$ $=\frac{\partial}{\partial s_{i}}u_{i}(w^{t},s^{t})+\mu_{i}^{t},$
$\text{ where }\mu_{i}^{t}$ $=\begin{cases}-\frac{\partial}{\partial s_{i}}u_{i}(w^{t},s^{t}),&\text{when either }s_{i}^{t}=0,\frac{\partial}{\partial s_{i}}u_{i}<0\\ &\text{or }s_{i}^{t}=s_{i}^{\max},\frac{\partial}{\partial s_{i}}u_{i}>0\\ 0.&\text{otherwise}\end{cases}$
@@
The mechanism is detailed out in Algorithm˜1. Our main result of this section is that under certain bounded derivative conditions, Algorithm˜1 always converges to a Nash equilibrium. We need a few matrices, defined as follows for every $w$ and $s$. For $i,j\in N$ and $k,\ell\in\{1,\ldots,m\}$,
$\begin{split}\textstyle&\textstyle G(w,s)_{ij}=\frac{\partial^{2}}{\partial s_{j}\partial s_{i}}u_{i}(w,s);\>\tilde{G}(w,s)_{k\ell}=\frac{1}{n}\sum_{i\in N}\frac{\partial^{2}}{\partial w_{\ell}\partial w_{k}}v_{i}(w,s);\\ \textstyle&\textstyle H(w,s)_{ik}=\frac{\partial^{2}}{\partial w_{k}\partial s_{i}}u_{i}(w,s);\>\tilde{H}(w,s)_{kj}=\sum_{i\in N}\frac{\partial^{2}}{\partial s_{j}\partial w_{k}}v_{i}(w,s).\end{split}$ (5)
@@
Consider the utility functions given by Equation˜1 where functions $v_{i},z_{i},i\in N$ are such that the following properties hold for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$ (the matrices below are as defined in Equation˜5).
@@
    We assume the following bounds:
    1. (a)
      $\forall i,j\in N,|G(w,s)_{ij}|\leqslant L$,
    2. (b)
      $\forall k,\ell\in\{1,\ldots,m\},|\tilde{G}(w,s)_{k\ell}|\leqslant\tilde{L}$,
    3. (c)
      $\|H(w,s)\|_{op}\leqslant P$, where $\|A\|_{op}:=\inf\{c\geqslant 0:\|Ax\|\leqslant c\|x\|,\forall x\}$,
    4. (d)
      $\|\tilde{H}(w,s)\|_{op}\leqslant\tilde{P}$.
Note that the definition of $G$ and $\tilde{G}$ do not imply the strong concavity of $u_{i}$ or $v_{i}$. Moreover, the above assumptions are standard and have appeared in the literature before (see (Murhekar et al., 2023)). We define the following expressions for a cleaner presentation.
$\begin{split}\textstyle W_{1}=\sqrt{1+\gamma^{2}n^{2}L^{2}-2\gamma\lambda}&\textstyle+\sqrt{\tilde{P}^{2}\gamma^{2}},\\ \textstyle W_{2}=\sqrt{1+\eta^{2}m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda}}&\textstyle+\sqrt{P^{2}\eta^{2}},\\ \textstyle E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2},&\textstyle\>\text{ where }w^{0}\in\mathbb{R}^{m},s^{0}\in\prod_{i\in N}S_{i}\\ \textstyle T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{\textstyle E}{\textstyle\epsilon}}\right)}}{{\left(\ln{\frac{\textstyle 1}{\textstyle W}}\right)}},&\textstyle\>\text{ where }W=\max\{W_{1},W_{2}\}.\end{split}$ (6)
We are now ready to present the main result of this section. Owing to the space constraints, all proofs of the technical results are presented in the Appendix.
@@
Under ˜1, UPBReD (Algorithm˜1) converges to a Nash equilibrium. Formally, for a given $\epsilon>0$, for every initial value $(w^{0},s^{0})$ of Algorithm˜1, the gradients $\|g(w^{T},s^{T},\mu^{T})\|<\epsilon$ and $\|\tilde{g}(w^{T},s^{T})\|<\epsilon$, for all $T\geqslant T_{0}(w^{0},s^{0})$, when the step sizes are chosen as follows:
$\gamma$ $<\min\left\{1,\frac{1}{\tilde{P}},\frac{2\lambda}{n^{2}L^{2}},\frac{\lambda-\tilde{P}}{n^{2}L^{2}-\tilde{P}^{2}}\right\},\text{ given }\lambda>\tilde{P}$,
$\eta$ $<\min\left\{1,\frac{1}{P},\frac{2\tilde{\lambda}}{m^{2}\tilde{L}^{2}},\frac{\tilde{\lambda}-P}{m^{2}\tilde{L}^{2}-P^{2}}\right\},\text{ given }\tilde{\lambda}>P.$
Proof.

### v4

@@ 3 Learning with arbitrary effective costs
In this section, we consider the effective cost of agent $i$, i.e., $z_{i}(s_{i},s_{-i})$, to be an arbitrary function of $s=(s_{i},s_{-i})$. We know that an alternative interpretation of NE (Definition˜1) is a strategy profile $(s_{i}^{*},s_{-i}^{*})$ where every agent’s best response to the strategies of the other players is its own strategy in that profile (see  (Maschler et al., 2020, e.g.)), i.e., $s_{i}^{*}\in\operatorname*{\arg\!\max}_{s_{i}\in S_{i}}\ u_{i}(w,s_{i},s_{-i}^{*}),\ \forall i\in N$. Hence, an algorithm that simultaneously updates all agents’ strategies with the best responses to the current strategies of the other players is called a best response dynamics of a strategic form game (see  (Fudenberg, 1991, e.g.) for a detailed description). In our problem, this approach cannot be directly employed since the center also needs to learn and update the model parameters $w$ as the agents choose their data contributions. We, therefore, propose a mechanism for FL that simultaneously updates both the agents’ strategies and the center’s choice of $w$. FL protocols iteratively use local gradients to update a global model, the mechanism we describe below captures strategic behaviour at every iteration by allowing agents to update their strategy (data set sizes used for gradient computation) at every realisation of the global model based on Equation˜1.
@@
In this mechanism, each agent $i$ starts with some initial choices of $s_{i}^{0}$ and shares that with the center. [^1] Note that, only the number of data points, $s_{i}$ is shared with the center and not the data, which is consistent with the principle of FL. The center initializes a $w^{0}$ and broadcasts that and the entire initial data-contribution vector $s^{0}$ to all the agents. In every subsequent iteration $t$, each agent $i$ locally computes two quantities: (i) its updated contribution $s_{i}^{t+1}$ by taking one gradient ascent step w.r.t. its own contribution $s_{i}$, and (ii) the local gradient of agent $i$’s valuation component of the social welfare w.r.t. $w$. Both are evaluated at the current values of $w^{t}$ and $s^{t}$ and sent back to the center. The center calculates an updated $w^{t+1}$ that averages (in spirit of the FedAvg algorithm (McMahan et al., 2017)) all the local gradients sent by the agents. The center then shares $w^{t+1}$ and $s^{t+1}$ with all the agents. Formally, the updates are given as follows.
$s^{t+1}=s^{t}+\gamma g(w^{t},s^{t},\mu^{t}),$ $\qquad w^{t+1}=w^{t}+\eta\tilde{g}(w^{t},s^{t}),$ (3)
[…]
@@
The mechanism is detailed out in Algorithm˜1. Our main result of this section is that under certain bounded derivative conditions, Algorithm˜1 always converges to a Nash equilibrium. We need a few matrices, defined as follows for every $w$ and $s$. For $i,j\in N$ and $k,\ell\in\{1,\ldots,m\},G(w,s)_{ij}=\frac{\partial^{2}}{\partial s_{j}\partial s_{i}}u_{i}(w,s);\tilde{G}(w,s)_{k\ell}=\frac{1}{n}\sum_{i\in N}\frac{\partial^{2}}{\partial w_{\ell}\partial w_{k}}v_{i}(w,s);H(w,s)_{ik}=\frac{\partial^{2}}{\partial w_{k}\partial s_{i}}u_{i}(w,s)$; $\tilde{H}(w,s)_{kj}=\frac{1}{n}\sum_{i\in N}\frac{\partial^{2}}{\partial s_{j}\partial w_{k}}v_{i}(w,s)$.
@@
Consider the utility functions given by Equation˜1 where functions $v_{i},z_{i},i\in N$ are such that the following properties hold for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$, $\forall i,j\in N,\forall k,\ell\in\{1,\ldots,m\}$ (the matrices below are as defined above).
@@
    We assume the following bounds: (i) $|G(w,s)_{ij}|\leqslant L$, (ii) $|\tilde{G}(w,s)_{k\ell}|\leqslant\tilde{L}$, (iii) $\|H(w,s)\|_{op}\leqslant P$, where $\|A\|_{op}:=\inf\{c\geqslant 0:\|Av\|\leqslant c\|v\|,\forall v\}$, (iv) $\|\tilde{H}(w,s)\|_{op}\leqslant\tilde{P}$.
Note that the definition of $G$ and $\tilde{G}$ do not imply the strong concavity of $u_{i}$ or $v_{i}$. Moreover, the above assumptions are standard and have appeared in the literature before (see Murhekar et al. (2023)). We define the following expressions for a cleaner presentation. $W_{1}=\sqrt{1+\gamma^{2}n^{2}L^{2}-2\gamma\lambda}\sqrt{\tilde{P}^{2}\gamma^{2}},W_{2}=\sqrt{1+\eta^{2}m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda}}+\sqrt{P^{2}\eta^{2}},E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2},T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{E}{\epsilon}}\right)}}{{\left(\ln{\frac{1}{W}}\right)}},$ where $w^{0}\in\mathbb{R}^{m},s^{0}\in\prod_{i\in N}S_{i}$ and $W=\max\{W_{1},W_{2}\}$. We are now ready to present the main result of this section. Full proofs of the technical results are presented in the Appendix.
@@
Under ˜1, UPBReD (Algorithm˜1) converges to a Nash equilibrium. Formally, for a given $\epsilon>0$, for every initial value $(w^{0},s^{0})$ of Algorithm˜1, the gradients $\|g(w^{T},s^{T},\mu^{T})\|<\epsilon$ and $\|\tilde{g}(w^{T},s^{T})\|<\epsilon$, for all $T\geqslant T_{0}(w^{0},s^{0})$, when the step sizes are chosen as follows: $\gamma<\min\left\{1,\frac{1}{\tilde{P}},\frac{2\lambda}{n^{2}L^{2}},\frac{\lambda-\tilde{P}}{n^{2}L^{2}-\tilde{P}^{2}}\right\},\text{ given }\lambda>\tilde{P}$, and $\eta<\min\left\{1,\frac{1}{P},\frac{2\tilde{\lambda}}{m^{2}\tilde{L}^{2}},\frac{\tilde{\lambda}-P}{m^{2}\tilde{L}^{2}-P^{2}}\right\},\text{ given }\tilde{\lambda}>P$.
1
Output: $w^{T}$
2
3for $t=0$ to T1T-1 do
4 Center broadcasts $w^{t},s^{t}$
5
6 foreach agent $i\in N$ in parallel do
7    $s^{t+1}_{i}=s^{t}_{i}+\gamma[g(w^{t},s^{t},\mu^{t})]_{i}$
9      Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
10
11
13
14
15return $w^{T}$
16
Algorithm 1 Updated Parameter Best Response Dynamics (UPBReD)
The theorem provides a guarantee of convergence of Algorithm˜1 for any arbitrary initial condition $(w^{0},s^{0})$. It needs to run for a minimum number of iterations $T_{0}(w^{0},s^{0})$, with appropriate parameters of $\gamma$ and $\eta$ that govern the gradient ascent rates of the agents’ and center’s objectives respectively. Note that for a better convergence, i.e., a smaller $\epsilon$, the algorithm needs to run longer. The parameter $W$ determines the contraction factor of the recurrence of $\|g(w^{t},s^{t})\|_{2}+\|\tilde{g}(w^{t},s^{t})\|_{2}$ and is chosen to be smaller than unity by choosing $\gamma$ and $\eta$ appropriately.
(a) $\tilde{P}=1$
(b) $\lambda=1.1$
Remark. The step sizes $\gamma$ and $\nu$ in our analysis scale inversely with $m$ and $n$, and certain restrictions on $\lambda$ and $\tilde{\lambda}$ may appear restrictive. In Section 4, we address these issues and propose a learning algorithm with step sizes independent of $m$ and $n$.
3.1 Application of Theorem˜1: effective costs
are only personal
Example 1 (NE different from socially optimal).
So, our objective in this paper is to allow monetary transfer among the agents so that we get the best of both worlds: (a) We achieve convergence to a Nash equilibrium of $(w^{\texttt{{OPT}}},s^{\max})$, where all agents to contribute their entire data and center learns the optimal parameter, and (b) the transfers are budget balanced, i.e., the center does not accumulate any money – it is used just to realign the agents’ utilities to reach the desired Nash equilibrium. We discuss this mechanism in Section˜4.
While agents strategically choose $s_{i}$ to maximize their utility, we assume truthful reporting of these choices. In UPBReD, agents report $s_{i}^{t}$ and $d_{i}^{t}$ in each round. Section˜A.2 shows an example where misreporting can improve utility. The next subsection introduces payments to ensure truthful reporting of $s_{i}^{t}$ and $d_{i}^{t}$.
3.2 Truthful elicitation with payment rules
So far, we have assumed that the agents report their type truthfully in every round, i.e. if an agent is training on $s_{i}^{t}$ amount of data at time $t$ the agent reports $\theta_{i}=(s_{i}^{t},d_{i}^{t})$ truthfully. However, an agent can misreport their strategy and corresponding gradient update to improve their utility as observed in Section˜A.2.
A well established goal in social choice theory is to choose an allocation that maximizes the social welfare (Green and Laffont, 1977). In our setting this maximizing allocation for a given type profile $\theta^{t}$ at round $t$ will be $a_{*}({\theta}^{t})=\operatorname*{\arg\!\max}_{w\in\mathbb{R}^{m}}{\sum_{i\in N}v_{i}(w,s^{t}})$
However, in our FL setting, we cannot compute this welfare maximizing allocation. Instead, we use $a_{w}^{t}$, which iteratively takes gradient ascent steps towards the welfare maximizing allocation as shown in Algorithm˜1. The initial allocation is chosen as $w^{0}$. And future allocations are computed iteratively as $a_{w}^{t}(\theta^{t})=w^{t}(s^{t},d^{t})=w^{t-1}+\frac{\eta}{n}\sum_{i\in N}d_{i}^{t}$. Due to this difference, we need an additional assumption for truthfulness. We assume the difference between the welfare at our computed allocation and that of the maximum at truthful reports is bounded by a constant.
Assumption 2.
Remark. Note that the above assumption is weaker than that of bounded valuations, that are standard when choosing allocation that are welfare maximizing. When valuations are non-negative and bounded above by some constant $M$, $\Gamma$ can be upper-bounded by $n.M$. Our mechanism remains stable for any $\Gamma^{\prime}\geq\Gamma=n.M$, ensuring that the center only needs a coarse estimate of the maximum valuation to satisfy ˜2.
Consider the following payment rule:
where, $\phi=0\ \text{if }\hat{v}^{t}_{i}=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t});\phi=\Gamma+(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t}))^{2}\>\text{ otherwise}$.
  1. 1.
  2. 2.
  3. 3.
    Agents share their type as $\hat{\theta}_{i}^{t}=(\hat{s}_{i}^{t},\hat{d}_{i}^{t})$ which can be different from their true type. The center computes an allocation using the reported types using $a_{w}^{t}$ consistent with Equation˜3.
  4. 4.
[…]
  5. 5.
With the given allocation rule $a_{w}^{t}$ and payment rule $p$ described in Equation˜4 for a true type and valuation profile $\theta^{t},v^{t}$, a reported type and valuation profile $\hat{\theta}^{t},\hat{v}^{t}$, the utility realised by agent $i$ at round $t$ is given by $u_{i}(\hat{\theta}^{t},\hat{v}^{t}|\theta^{t},v^{t})=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),s^{t})-z_{i}(s_{i}^{t},\hat{s}_{-i}^{t})+p_{i}^{*}(\hat{\theta}^{t},\hat{v}_{i}^{t})$. The algorithm for T-UPBReD is detailed in Algorithm˜3 of the Appendix.
With this mechanism, we can show the following result, the proof of which is deferred to Section˜A.2.
Theorem 2.
Under ˜2, T-UPBReD is EPIC in every round.
In Section˜3.1, we present an example where the NE is not socially optimal, motivating the need for a method to achieve the social optimum as a NE, which we address in the following section.
4 Welfare maximization at Nash equilibrium
In this section, we consider the scenario where the center can make a payment $p_{i}(s_{i},s_{-i})$ to agent $i$ to alter its utility function. Therefore, the effective cost of agent $i$ becomes $z_{i}(s_{i},s_{-i})=c_{i}(s_{i})-p_{i}(s_{i},s_{-i})$. Note that, the payment can be either positive or negative, which determines whether the agent is subsidized or taxed respectively. It is reasonable to expect that the data contributors of the FL process are subsidized for their data contribution while the data consumers are charged payment for obtaining the learned parameter. We assume that the derivatives of the valuation $v_{i}$ and cost function $c_{i}$ are bounded. We choose the following payment function:
$p_{i}(s)=\beta(s_{i}-\frac{1}{n-1}\sum_{j\neq i}s_{j}),$ (5)
where $\beta$ is a parameter of choice. Note that this payment mechanism is budget balanced by design, since $\sum_{i\in N}p_{i}(s_{i},s_{-i})=0,\forall s_{i}\in S_{i},s_{-i}\in S_{-i}$. With payment, the utility of agent $i$ becomes $u_{i}(w,s)=v_{i}(w,s)-c_{i}(s_{i})+p_{i}(s).$ In this section, we are also interested in the quality of the NE in terms of social welfare and want to reach the desired NE where $s_{i}^{*}=s_{i}^{\max},\forall i\in N$ and $w^{*}=w^{\texttt{{OPT}}}$.
4.1 Convergence to the NE $(w^{\textnormal{OPT}},s^{\max})$
Assumption 3.
Note that the above assumptions are pretty mild. Previous works assume stronger assumptions (like convexity) on $v_{i}$ and $c_{i}$ (see Murhekar et al. (2023)). We show that the above bounds on the derivatives of $v_{i}$ and $c_{i}$ are sufficient to claim our result.
Lemma 1 (Increasing utility).
Suppose Assumption 3 holds and we use the payment function given by Equation˜5. Then, the utility of every agent is increasing in its own data contribution, i.e., $\frac{\partial}{\partial s_{i}}u_{i}(w,s_{i},s_{-i})>0,\forall s_{i}\in(0,s_{i}^{\max}),\forall i\in N,\text{ and }\forall w$ provided $\beta>\zeta+\tau$.
Definition 4 ($M$-smoothness).
Definition 5 ($\nu$-strictly convex).
1
2
Output: $w^{T}$
3 Center broadcasts $w^{0},s^{0}$, set $\tau=0$
4while $s\neq s^{\max}$ do
5 Center broadcasts $s^{\tau}$
6 for agent $i\in N$ in parallel do
7    $s_{i}^{\tau+1}=s_{i}^{\tau}+\gamma[g(w^{0},s^{\tau},\mu^{\tau})]_{i}$, send $s_{i}^{\tau+1}$ to center
8
9 $\tau=\tau+1$
10for $t=0$ to $T$ do
11 Center broadcasts $w^{t}$
12 for agent $i\in N$ in parallel do
13      Compute $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{\max},s_{-i}^{\max})$
14      Send $d_{i}^{t+1}$ to the center
15
17return $w^{T}$
Algorithm 2 2P-UPBReD
We propose a two-phase algorithm given by Algorithm˜2. The algorithm, in the first phase, incentivizes the agents to contribute $s^{\max}$, and in the second phase, converges to $w^{\texttt{{OPT}}}$. Note that in the first phase of Algorithm˜2, every agent runs a gradient ascent step. Thanks to the increasing utility property (Lemma 1), we have an increasing sequence of $s_{i}^{t}$ for all $i\in N$. Moreover, from the definition of $g(w^{0},s^{t},\mu^{t})]_{i}$ it is ensured that $s_{i}^{\max}$ is the fixed point of this gradient ascent update. Hence after a finite number of iterations, every agent reaches $s_{i}^{\max}$. Note that in this phase the model parameter $w^{0}$ remains unchanged.
In the second phase of the algorithm, we update the model parameter $w^{t}$. Note that since all the agents contribute the maximum amount of data they own, this phase is simply pure FL (without incentive design). As such, each agent now computes the local gradient $d_{i}^{t+1}$ and sends it to the center in each round.
We now provide the convergence guarantees of Algorithm˜2. Before that, let us discuss the necessary assumptions.
Assumption 4.
The smoothness and strong convexity assumptions have featured in several previous papers on FL Karimireddy et al. (2020b); Yin et al. (2018). The main result of this section is as follows.
Theorem 3 (2P-UPBReD convergence to the optimal NE).
Suppose Assumptions  3 and 4 hold and we consider the utility function with payment scheme given in Equation˜5 with $\beta>\zeta+\tau$. Then, for every $w^{0}\in\mathbb{R}^{m}$ and $s^{0}\in\prod_{i\in N}S_{i}$, 2P-UPBReD (Algorithm˜2) converges to the Nash equilibrium $(w^{\texttt{{OPT}}},s^{\max})$ and is budget balanced. In particular, for any given $\epsilon>0$, we have $\frac{1}{n}\sum_{i\in N}v_{i}(w^{\texttt{{OPT}}},s^{\max})-\frac{1}{n}\sum_{i\in N}v_{i}(w^{T},s^{\max})<\epsilon$, and $s^{T}=s^{\max}$ in $T=\kappa+T_{0}$ iterations provided we choose $\gamma=c$ (an universal constant) and $\eta=1/M$ with
$\kappa\geqslant\max_{i}\left\{\frac{(s^{\max}_{i}-s^{0}_{i})}{c\,\Delta}\right\},$ (6)
and $T_{0}>\left(\ln{\frac{f(w^{0},s^{\max})-f(w^{\texttt{{OPT}}},s^{\max})}{\epsilon}}\right)\Big/\left(\ln\frac{1}{{1-\frac{\nu}{M}}}\right)$ where, $f(w,s):=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s)$ and $\Delta=\beta-\tau-\zeta$.
Remark (Strong convexity and smoothness). The convergence rate is $\mathcal{O}(\ln{\nicefrac{{1}}{{\epsilon}}})$ which is the same as that of Algorithm˜1. However, Algorithm˜2 is cleaner in terms of the phases of convergence, the choices of the step sizes $\gamma,\eta$. The structural assumptions like strong convexity and smoothness may be relaxed at the expense of weaker rates. For example, we observe convergence rates of $\mathcal{O}(\nicefrac{{1}}{{\sqrt{\epsilon}}})$ for non-convex smooth functions and $\mathcal{O}(\nicefrac{{1}}{{\epsilon}})$ for convex smooth functions.
Remark ($\gamma$ and rationality). Phase 1 of Algorithm˜2 achieves the optimal contribution $s^{\max}$ without updating the learning parameter $w$. The speed of convergence of this phase depends on the parameter $\gamma=c$, which can be interpreted as the minimum rationality level of the society. For instance, if all agents are sufficiently rational, i.e., $c$ is large, they may converge to $s^{\max}$ in a single iteration. However, the algorithm converges even if agents are boundedly rational. In Phase 2 of the algorithm, the agents stop updating their contributions and only update the gradients $d_{i}$ and the center accumulates them to update $w^{t}$.
Thanks to the smoothness and strong convexity properties of $f(\cdot,s^{\max})$, we can show that running Algorithm˜2 guarantees that the model iterate $w^{T}$ converges to $w^{\texttt{{OPT}}}$ at an exponential speed as well.
Corollary 1 (Iterate Convergence).
With the same setup as above and $\gamma=c$ (universal constant), $\eta=\frac{2}{M+\nu}$, we obtain $\|w^{T}-w^{\texttt{{OPT}}}\|_{2}<\epsilon,\text{ and }s^{T}=s^{\max}$, where $T=\kappa+\tilde{T}_{0}$, with the same $\kappa$ as in Equation˜6 and $\tilde{T}_{0}>\left(\ln{\frac{\|w^{0}-w^{\texttt{{OPT}}}\|_{2}}{\epsilon}}\right)/\left(\ln\frac{{1+\frac{\nu}{M}}}{{1-\frac{\nu}{M}}}\right)$.
Remark (Truthfulness). An agent’s true type is $(s_{i}^{max},d_{i}^{t})$ in every iteration of phase $2$. We can use the payment scheme $p_{i}^{*}$ as defined in Equation˜4, with the necessary assumptions, in addition to the $z_{i}(s_{i},s_{-i})$ defined in this section, making the mechanism truthful with the trade-off that it will no longer be budget-balanced. Green and Laffont (1979) show that it is impossible to achieve BB, truthfulness and optimality together.
Remark (Robustness to adversarial behavior). In 2P-UPBReD, we have assumed that the agents report their gradients truthfully in phase $2$. If the agents misreport by sending arbitrary gradients not computed on any data, we have discussed one option of enforcing additional payment in the previous remark that compromises budget balance. If we do not wish to enforce additional payments, we can use some established methods in literature that are robust to adversarial agents (Yin et al., 2018; Karimireddy et al., 2020a; Ghosh et al., 2021). We describe one such method called trimmed mean  (Yin et al., 2018), in Section˜B.4 and implement 2P-UPBReD with trimmed mean in the following experiments
In Section˜B.4, we describe the trimmed mean algorithm for 2P-UPBReD designed to be robust in the presence of adversarial agents. In this section, we demonstrate this robustness on the FEMNIST and CIFAR-10 dataset. We vary the adversarial fraction $\alpha\in\{0.025,0.05,0.075,0.1,0.125,0.15,0.175,0.2\}$, selecting suitable trimming parameters. Figure˜2 shows the ratio of social welfare with and without adversaries for UPBReD and 2P-UPBReD. While both suffer welfare loss under adversaries, applying the trimmed mean algorithm (Algorithm˜4) keeps the welfare ratio close to $1$ for trimmed mean 2P-UPBReD.
(a) CIFAR-10
(b) FEMNIST
5 Experiments
Our results in Sections˜3 and 4 guarantee convergence of Algorithms˜1 and 2 to a pure Nash equilibrium under the given assumptions. However, it is important to ask how these FL algorithms behave in practice with real datasets under more general setups. We capture results in these realistic settings in our experiments. We train models on the CIFAR-10  Krizhevsky (2009), FEMNIST, and Twitter datasets Caldas et al. (2019) and compare the social welfare of UPBReD, 2P-UPBReD, and FedAvg. We also consider another algorithm FedAvgStrategic as a baseline. In this algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg (details in Section˜C.1). The mechanisms/algorithms proposed in Table˜1, which are closest to this paper, do not use valuation functions that consider simultaneous learning and strategy updates and are therefore not meaningful to compare against. The cost function used is linear $c_{i}(s_{i})=c_{i}\cdot s_{i}$, where $c_{i}$ is sampled from $U[0,1]$ for each agent. We choose our valuation function to be $v_{i}(w^{k},s^{k})=r_{i}-\mathcal{L}_{i}(w^{k},s^{k})$ where the $\mathcal{L}_{i}(w^{k},s^{k})$ is the loss evaluated on agents $i$’s test dataset at the $k$-th round of training. The social welfare is given by $\sum_{i\in N}v_{i}(w^{k},s^{k})$. We perform strategy updates using the numerical derivative as follows: $s_{i}^{k+1}=s_{i}^{k}-\frac{L_{i}(w_{i}^{k})-L_{i}(w^{k-1})}{s_{i}^{k}-s_{i}^{k-1}}-c_{i}+\beta$.
The model parameters obtained by agent $i$ after performing one training step locally on the model initialized with $w^{k-1}$ is denoted by $w_{i}^{k}$. We use the cross-entropy loss function for $L_{i}$} on agent i’s dataset. All the experiments are performed on NVIDIA RTX A5000 and NVIDIA RTX A6000 with a 24GB and 48GB core respectively. The experiments required to obtain the results we report take several days to run. For comparing social welfare against $\beta$ and number of agents, we randomize over 10 runs, sampling an integer $s^{0}_{i}$ from $U[\frac{s^{\max}_{i}}{3},\frac{2s^{\max}_{i}}{3}]$ and cost from $U[0,1]$. We plot the average value of time taken across $10$ runs.
Dataset specifics. For CIFAR-10, the data is distributed equally among 100 agents by sampling the data uniformly at random. We train a CNN with 1,250,858 parameters for 100 rounds. Hyperparameters: $\gamma=0.5,\ \beta=2,\ \eta=0.001$, optimizer=Adam.
For FEMNIST, we use $20\%$ of the dataset that generates users with data that is not identically distributed as per the codebase provided by Caldas et al. (2019). We train a CNN with $6,603,710$ parameters for $100$ rounds. Hyperparameters: $\gamma=0.1,\ \beta=2,\ \eta=0.001$, optimizer=Adam.
For Twitter, we use $15\%$ of the dataset that generates users with data that is identically distributed. We further group $1000$ agents together to create agents with larger data shares. We use an LSTM network to train for $50$ rounds. Hyperparameters: $\gamma=0.5,\ \beta=2,\ \eta=0.00001$, optimizer=Adam.
All results are summarized in Figure˜3. We provide more details of the model architecture for all models in Section˜C.2.
(a) CIFAR-10
(b) FEMNIST
(c) Twitter
Figure 3: The three rows in this figure correspond to the experiments with the datasets CIFAR-10, FEMNIST, and Twitter, respectively. The y-axis in every plot is the social welfare. The x-axis corresponds to the number of agents in the first plot, the choice of $\beta$ in the second, and the time taken (in seconds) in the third plot, respectively, in every row.
We observe that 2P-UPBReD and FedAvg have the same social welfare performance w.r.t. the number of agents (first column), and running times are nearly the same for all the datasets (third column). 2P-UPBReD outperforms UPBReD in nearly all the runs with respect to social welfare. We also observe that 2P-UPBReD converges to the maximum contribution of the agents for values of $\beta$ that are lower than the theoretically predicted ones. We also note that the data contributors get paid and data consumers pay in 2P-UPBReD and their payments are monotone non-decreasing with data contribution (Figure˜4). The utilities vary because the dataset on which the model is tested to obtain the accuracy varies for the agents.
We leverage the partitioners provided by the Flower framework to generate multiple data splits of the CIFAR-10 dataset. We evaluate the performance of the algorithms under three data-partitioning regimes: an IID split, a Dirichlet split, and a Pathological split. The resulting distribution for the Dirichlet and Pathological split is visualised in Figure˜6. For the Dirichlet split, the concentration parameter is set to $\alpha=0.1$ with self-balancing enabled. For the Pathological split, the class assignment mode is set to random, and each partition contains $3$ classes. Hyperparameter details: $\beta=2,\gamma=1,\eta=0.0001$,optimizer=Adam. The results are summarized in Figure˜5, averaged across $5$ runs.
In Section˜B.4, we describe the trimmed mean algorithm for 2P-UPBReD designed to be robust in the presence of adversarial agents. In LABEL:app:experiments_adversarial we show that 2P-UPBReD with trimmed mean is robust to adversarial agents.
(a) IID Split
(b) Dirichlet Split
(c) Pathological Split
Figure 5: Social welfare vs. number of clients for different CIFAR-10 dataset splits using the Flower framework.
(a) Pathological partitioner (33 classes per client).
(b) Dirichlet partitioner ($\alpha=0.1$).
Figure 6: Label distributions for different data partitioning strategies using Flower with $10$ agents. The x-axis represent the client id and the y-axis represents the number of data points of the given label.
6 Conclusions and future work
Appendix A Appendix to Section˜3
A.1 Proof of Theorem˜1